A 1030. kg automobile is pulled by a horizontal tow line with a net force of 834 N. What is the acceleration of the auto?
1
Expert's answer
2013-04-30T09:10:25-0400
Due to the 2nd Newton Law, F = m * a.
The mass is m = 1030 kg, the net force is F = 834 N. Thus the acceleration
a = F / m = 834 N / 1030 kg = 0.8097 m/s^2.

Answer: about 0.81 m/s^2